# SynthesizRR: Generating Diverse Datasets with Retrieval Augmentation

### v1

@@ 6 Results: Student distillation
Analysis of student distillation performance. Why does SynthesizRR improve classification dataset synthesis? We take a closer look at the few-shot generation strategies from a student training perspective.
@@
We first look at the label preservation accuracy, obtained from an oracle DeBERTa-v3L model tuned on the full Train dataset (Appendix H). We would expect that better label preservation means a higher-fidelity training dataset. However, Table 7 shows that FewGen datasets have very high label preservation in spite of their lower test performance. Especially on multiclass tasks (AG., ToI, Cat.), FewGen shows the highest label preservation (exceeding Gold) but this does not translate into improved student performance.
@@
To understand this, we conduct a deeper analysis of the student training dynamics on multiclass datasets. We train a DistilBERT student for 6 epochs and plot the corresponding data-maps Swayamdipta et al. (2020). As noted in §4, label preservation can be exploited by only generating easy to classify examples. From the Category data maps in Figure 5, it is clear that FewGen generations tend to cluster around easy-to-learn examples (high confidence and low variability). SynthesizRR contains more ambiguous examples (high variability) which we hypothesize are essential to learning the nuances between classes. To validate this, we filter out the least ambiguous examples (bottom 17% variability) and retrain the students. In Table 8 we find that FewGen performance degrades, whereas synthesizer improves (giving us new best performances, despite using only 83% of rows). We conclude that SynthesizRR generates more ambiguous examples, and this helps establish better class-separability in multi-class data sets.
@@
SynthzRR $87.56$ 72.8 $74.79$ $69.41$ 90.7 $99.32$
SynthzRR $87.36$ $65.85$ $73.24$ $73.24$ $77.35$ 99.7
Table 7: Few-shot label-preservation accuracy ($\uparrow$) using tuned oracle DeBERTa-v3L model. Gold row is accuracy on 20% validation split. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
@@
Table 8: Test Accuracy ()($\uparrow$)( ↑ ) after keeping 83% most-ambiguous examples. We report improvements compared to Table 6. Notation: *32-shot; 3-shot Retr-ICL; 32-shot No-Retr-ICL.
7 Comparison to additional baselines
Method Teacher / Retriever Data Size Yelp AG. IMDb SST-2
Gold - 6k 94.5 90.8 91.3 88.2
ZeroGen GPT2-XL 6k
$\otimes$
$\otimes$
[…]
82.4
(Reported) 200k
$\otimes$
$\otimes$
84.3 87.3
ProGen GPT2-XL 6k
[…]
(Reported) 100k 89.4 80.8 84.1 87.2
SunGen GPT2-XL 6k
[…]
84.9
[…]
(Reported) 200k 89.2 80.5 89.4 89.5
ReGen BERT 6k
[…]
82.7
[…]
(Reported) 20k 89.0 85.0 87.8 85.3
S3 GPT3.5-T 6k
$\otimes$
$\otimes$
87.1
$\otimes$
(Reported) 21.2k
$\otimes$
$\otimes$
89.0
$\otimes$
AttrPmt GPT3.5-T 6k 89.3 79.8
[…]
80.8
(Reported) 6k 90.6 82.7 85.9 83.6
Zero shot
SynthzRR LLaMa2 $+$ Contr. 6k 85.6 85.3 82.9 80.2
SynthzRR ClaudeV1 $+$ Contr. 6k 87.5 85.6 83.6 82.5
SynthzRR LLaMa2 $+$ BM25 6k 83.2 84.3 74.1 84.4
SynthzRR ClaudeV1 $+$ BM25 6k 86.7 84.2 79.1 82.6
3-shot Retr-ICL
SynthzRR LLaMa2 $+$ Contr. 6k 89.9 84.6 84.8 83.8
SynthzRR ClaudeV1 $+$ Contr. 6k 91.3 86.0 86.3 80.6
SynthzRR LLaMa2 $+$ BM25 6k 90.7 84.3 84.7 84.4
SynthzRR ClaudeV1 $+$ BM25 6k 89.4 84.1 81.6 82.3
Table 9: Test accuracy of SynthesizRR Retr-ICL against baselines after fine-tuning a DistilBERT student. For each released dataset, we subsample to 6k examples to make a fair comparison. We train 5 student models and report the mean accuracy (std. dev. was $\leq 2.0$ in all cases, thus we avoid reporting it for brevity). We report the DistilBERT accuracy from the original paper (if available) in italics. For each dataset, we bold the best result from the fair evaluation. Tasks not evaluated by previous authors are denoted by
$\otimes$
, while those evaluated without dataset release are marked as
[…]
. GPT3.5-T is gpt-3.5-turbo.
Here, we explore how SynthesizRR directly compares to prior work on synthesis of text classification datasets. We compare against six prior approaches:
ZeroGen Ye et al. (2022a): a simple approach which uses zero-shot generation from GPT2-XL under different prompting schemes. For certain tasks like IMDb and SST-2, the authors occasionally seed the prompt with a previously-generated movie name (described subsequently).
ProGen Ye et al. (2022b): generates a small synthetic dataset, then uses a student model along with an influence function to highlight high-quality generations. These are used as in-context examples for another round of generation.
SunGen Gao et al. (2023): uses the ZeroGen strategy to generate a large synthetic dataset (200k rows). Then, uses a custom bi-level optimization algorithm (involving the student model) to determine instance-weights of each synthetic example.
ReGen Yu et al. (2023b): performs multi-round filtering of retrieved results using 2 BERT models; one trained for retrieval, and one classifier. Use consistency between these models to filter noisy data.
S3 Wang et al. (2023a): Constructs a “seed dataset” (different from ours) and trains a student model. Then, extrapolate errors using an LLM and synthesizes additional data. We combine this with the seed data and repeat the process.
AttrPrompt Yu et al. (2023a): a method focused on improving diversity and unbiasedness of generated datasets. Prompts a powerful LLM like GPT3.5-Turbo with different attributes, each along different dimensions. Attributes are extracted from a human-in-the-loop analysis of task using GPT3.5-Turbo.
We note that zero shot generation and few-shot generation has already been compared in Table 6, so we do not include it here.
We benchmark four classification tasks which are popular in prior work: Yelp Polarity (Zhang et al., 2015), IMDb Maas et al. (2011), SST-2 Socher et al. (2013) and AG News (Zhang et al., 2015). The first three tasks are binary sentiment analysis, whereas the last is multi-class topic classification.
For the content sourcing stage of SynthesizRR, we retrieve documents from the following corpora:
  • Yelp Polarity: to generate positive and negative reviews about businesses, we retrieve from the Yelp Dataset corpus Yelp (2022) which contains details of 150k businesses. Each document contains the the name, address, and various tags and boolean attributes for a single business.
  • IMDb and SST-2: to generate movie reviews, we retrieve from the CMU Movie Summary corpus Bamman et al. (2013), which contains 42k plot summaries. Each document is a plot summary for a single movie.
  • AG News: we use RealNews/Dominant from Table 2.
We note that the CMU Movie Summary and Yelp Dataset corpora are between one to three orders of magnitude smaller than corpora in Table 2, thus it helps us evaluate SynthesizRR with smaller corpora. We include two different retrieval approaches: Contriever (semantic match) and BM25 (lexical match).
Following Yu et al. (2023a), we generate 6,000 rows using the SynthesizRR Retr-ICL variant described in Algorithm 1. We generate an equal number of examples for each class. We calculate intrinsic metrics Self-BLEU-5, MAUVE and Entity Entropy as described in §4.
Prior work often generates much larger datasets (20k to 200k examples) and uses different student model hyperparameters. Intrinsic evaluations of dataset quality are also seldom reported. This makes it difficult to fairly compare results from previous papers. Thus, we reproduce results ourselves by using synthetic datasets released by the authors, after subsampling to 6,000 examples. We report distillation results in Table 9 and intrinsic metrics in Table 10.
Method Teacher Retriever Self-BLEU-5 ()($\downarrow$)( ↓ ) Entity Entropy ()($\uparrow$)( ↑ ) Mauve ()($\uparrow$)( ↑ )
(Dataset) LM Yelp AG. IMDb SST-2 Yelp AG. IMDb SST-2 Yelp AG. IMDb SST-2
Gold - - 27.5 17.1 27.9 35.5 7.1 6.6 7.5 3.2 - - - -
ZeroGen GPT2-XL -
$\otimes$
$\otimes$
[…]
15.3
$\otimes$
$\otimes$
[…]
4.9
$\otimes$
$\otimes$
[…]
52.0
ProGen GPT2-XL -
[…]
SunGen GPT2-XL -
[…]
15.4
[…]
4.9
[…]
68.7
[…]
ReGen - BERT
[…]
56.5
[…]
8.1
[…]
68.1
[…]
S3 GPT3.5-T -
$\otimes$
$\otimes$
62.2
$\otimes$
$\otimes$
$\otimes$
5.7
$\otimes$
$\otimes$
$\otimes$
62.0
$\otimes$
AttrPmt GPT3.5-T - 70.3 39.8
[…]
71.5 3.4 6.0
[…]
3.4 55.9 52.8
[…]
50.0
Zero shot
SynthzRR LLaMa2 Contr. 70.2 29.3 66.3 41.9 5.4 7.1 5.7 4.5 67.9 89.5 58.5 50.0
SynthzRR ClaudeV1 Contr. 70.3 31.5 51.5 45.3 5.5 6.6 5.3 4.8 54.0 94.2 55.9 50.0
SynthzRR LLaMa2 BM25 70.6 28.7 62.2 36.5 4.7 7.0 5.6 5.1 64.7 90.3 60.5 50.0
SynthzRR ClaudeV1 BM25 72.9 30.9 50.4 36.9 4.5 6.5 5.1 5.4 54.6 90.8 53.2 50.0
3-shot Retr-ICL
SynthzRR LLaMa2 Contr. 59.5 34.2 62.9 26.3 4.9 7.2 5.7 3.8 84.9 92.6 72.6 50.0
SynthzRR ClaudeV1 Contr. 58.5 23.7 38.0 24.6 4.9 6.7 5.9 4.3 70.3 95.8 58.0 50.0
SynthzRR LLaMa2 BM25 66.6 32.0 59.7 25.3 5.1 7.2 5.6 4.8 81.9 92.5 78.7 50.0
SynthzRR ClaudeV1 BM25 59.9 24.6 41.9 26.8 5.0 6.7 5.4 4.9 63.3 96.0 58.5 50.0
Table 10: Intrinsic evaluations of synthetic datasets released by prior work. We subsample all to 6k examples before computing metrics as described in §4. Within each dataset, we bold the best result. Tasks not evaluated by previous authors are denoted by
$\otimes$
, while those evaluated without dataset release are marked as
[…]
. GPT3.5-T is gpt-3.5-turbo.
RQ: How does SynthesizRR compare to existing approaches in terms of distilled student accuracy?
We measure accuracy of a DistilBERT Sanh et al. (2019) student model. DistilBERT is a standard model used in prior dataset synthesis work Yu et al. (2023a); Ye et al. (2022a); Gao et al. (2023); Wang et al. (2023a); Ye et al. (2022b). We set the same model hyperparams as Yu et al. (2023a), i.e. training for 5 epochs using using Adam optimizer with batch_size=32 , lr=2e-5 , weight_decay=1e-4 and epsilon=1e-6 . We do not use cartography-based filtering for SynthesizRR to make a fair comparison. In addition to our results on 6k examples, we report original numbers from prior work (which might be on much larger datasets).
The results in Table 9 show that methods which rely on relatively weak LLM teachers like GPT2-XL Radford et al. (2019) can perform well on sentiment and topic tasks like IMDb, but require a very high data cost (15-30x more synthetic data than SynthesizRR). When scaled down to 6k rows, the performance of methods like ZeroGen and SunGen deteriorates.
Approaches which use strong instruction-following LLMs like AttrPrompt, S3, and SynthesizRR can achieve similar or better performance with much smaller datasets. These methods create high-quality datasets rather than modify the student modeling process, as is done by SunGen and ProGen. SunGen performs an iterative bi-level optimization over the ZeroGen datasets, jointly learning instance-weights and improving the student. ProGen uses an iterative filtering approach by training a student model and filtering partial generations based on a noise-robust, to curate an in-context dataset. We hypothesize these additions of the student model into the synthesis process also impact the final classification accuracy, as the dataset becomes specialized to the particular choice of student and its hyperparams. As we use a standard student-distillation setup here, datasets from these approaches may not perform as well.
More complex prompting techniques like Chain-of-Thought (Wei et al., 2022) used by S3 can indeed improve the task-inversion step, though this requires much higher API costs due to longer output lengths. Chain-of-Thought prompting thus seems like a promising approach by which SynthesizRR’s task-inversion step could be augmented.
In Table 10, we measure the entity diversity (Entity Entropy), lexical diversity (Self-BLEU), and similarity to Gold texts (MAUVE) compared to prior approaches. Among prior approaches, only AttrPrompt (Yu et al., 2023a, Appendix E) attempts to improve diversity of the generated text, by templatizing the task inversion instruction with attributes such as style, topic, length:min-words and more. ReGen is the only synthesis approach which uses content sourcing (but not task inversion). We thus consider these the two most relevant baselines for SynthesizRR.
We see that both ReGen and SynthesizRR achieve very high entity entropy compared to AttrPrompt, thus underscoring the importance of a content sourcing step. Unlike SynthesizRR, ReGen uses only retrieval, thus it has no explicit task-inversion step to make the contexts similar to Gold texts. Thus, we observe that ReGen suffers in terms of lexical diversity, Mauve and student accuracy, compared to SynthesizRR.
On the other hand, Chain-of-Thought prompting (S3), despite generating a strong classification dataset as we saw in Table 9, suffers from the lack of lexical diversity and similarity to Gold texts. This is also seen in AttrPrompt and previously in FewGen. This lends evidence to the hypothesis that synthesis without content sourcing tends to produce datasets with lower diversity, which cannot be overcome by complex prompting strategies alone.
Finally, we observe that ZeroGen shows very high diversity on SST-2, a task which involves generating partial movie reviews having positive or negative sentiment. However, as mentioned in Ye et al. (2022a, Section 4.6), ZeroGen is not simply zero-shot generation; the authors first generate a movie using the prompt Movie: and then insert the generated movie name into the prompt template The movie review in positive sentiment for movie "<Movie>" is: to generate an SST-2 example. We posit that the generated movie fulfils the same purpose as a retrieved context in SynthesizRR. Similarly high diversity is observed by SunGen which starts with ZeroGen-generated dataset and learns instance-weights (Gao et al., 2023, Section 2.2). Such synthetic content sourcing could be a fruitful direction for future work, to overcome the need for a retrieval corpus in SynthesizRR.
8 Related Work
Dataset synthesis using LLMs. Using LLMs to perform task inversion for dataset synthesis has been studied previously. Most use GPT-2XL without fine-tuning (Ye et al., 2022b, a; Gao et al., 2023; Meng et al., 2022; Schick and Schütze, 2021; Jung et al., 2023). Only recent works have considering large teacher LLMs such as GPT-3 (West et al., 2022; Honovich et al., 2023; Wang et al., 2023b) and chat-tuned LLMs such as gpt-3.5-turbo (Yu et al., 2023a; Yehudai et al., 2024b; Wang et al., 2023a).
For the generation of text classification datasets, class-conditioned prompting is key. Prior approaches have investigated zero-shot (Ye et al., 2022a), iterative few-shot prompting (Ye et al., 2022b), or synthesis using seq2seq LLMs fine-tuned on a curated dataset (Lee et al., 2021), but only recently has the importance of varying the prompt been established. Yu et al. (2023a) use prompts which explicitly attribute variations (such as subtopics and brands) and find this obtains greater diversity. We explore adding retrieval contexts to the prompt as the source of diversity and knowledge-grounding, and find it can significantly improve representation of tail entities. This avoids issues related to memorization of parametric knowledge (Roberts et al., 2020; Kandpal et al., 2023). We note that distillation for generation tasks like paraphrasing and summarization (Jung et al., 2023) has been demonstrated using off-the-shelf critic models; instead, our work focuses on classification and does not need a filtering step to be effective (though, as we show in §6, it does benefit from such steps in the multi-class case).
Retrieval-augmented generation. Our approach has many of the characteristics of in-context retrieval-augmented generation (RAG) (Lewis et al., 2020; Ram et al., 2023; Huang et al., 2023; Izacard et al., 2023). Previous studies show how RAG bypasses numerous problems associated with generating solely from parametric memory, i.e. heightened bias towards “head” entities (Mallen et al., 2023), lower lexical diversity (Holtzman et al., 2019; Jentzsch and Kersting, 2023), and hallucinated and confabulated information (Zhang et al., 2023). To our best knowledge, using retrieval-augmented generation for synthesis of classification tasks has not been explored at the instance level. Yu et al. (2023b) studied the retrieval-only setting for creation of topic and sentiment datasets; as described in §section 2 we consider tasks such as Hyperpartisan, Humor, and ToI Headlines, which pose additional challenges. Viswanathan et al. (2023) perform dataset-level retrieval and not instance-level retrieval, additionally selecting the student model based on the task whereas we focus on a fixed student. Finally, note that our application looks very different from other RAG applications like question answering (Asai et al., 2023) or counterfactual data augmentation (Paranjape et al., 2022); we generate diverse outputs only loosely connected to the input query.
Diversity in dataset synthesis. An important distinction between RAG and SynthesizRR is the desire for diverse generations. RAG has so far been used primarily for knowledge-based tasks such as open-domain question-answering, fact-checking and few-shot classification (Asai et al., 2023). These tasks face the issue of multiple valid surface-forms, which only change the wording of the generated output but not its meaning, making evaluation difficult (Holtzman et al., 2021). Greedy sampling is used most often.
On the other hand, dataset synthesis is emboldened by diverse surface-forms, as this produces diverse covariates linked to the same target, which intuitively aids generalization; we thus use top-p sampling (Holtzman et al., 2019). Another major difference from RAG is that we retain the long-tail of retrieved results, as this allows us to create a diverse grounded dataset.
9 Conclusion
Limitations
Appendix A Risks
Although the main goal of our work is to improve text classification, our use of LLMs to generate examples does carry some conceptual risks. By generating news articles to train classifiers on, we run the risk of generating fake news and other harmful content. However, we believe this risk is mitigated by the fact that the final outcome of our system is a classifier: classification models have relatively constrained failure modes (misclassification) compared to text generation models that can mislead users. Furthermore, we do not believe our approach uniquely advances the generation of fake news; our advances are largely orthogonal to the technology that brings such risks.
Appendix B Dense vs sparse retrieval in SynthesizRR
Retriever AG. Hyp. ToI Cat. Hum. Pol. Avg.
(Size) ($8$K) ($4$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold 91.0 93.2 82.5 81.5 93.1 95.3 89.43
LLaMa2 Zero shot
Contr. 83.5 69.8 74.5 68.9 82.5 84.7 77.32
BM25 83.2 74.2 70.7 57.6 78.5 85.4 74.93
ClaudeV1 Zero shot
Contr. 83.9 72.3 71.8 66.8 62.1 88.7 74.29
BM25 83.2 57.2 69.8 53.7 73.9 91.8 71.60
LLaMa2 3-shot Retr-ICL
Contr. 83.0 78.5 73.3 72.4 90.2 91.0 81.38
BM25 82.1 77.9 71.9 65.4 87.5 87.4 78.69
ClaudeV1 3-shot Retr-ICL
Contr. 83.7 72.3 72.8 65.4 83.4 91.3 78.16
BM25 83.0 73.5 70.0 52.4 82.4 90.7 75.34
Table 11: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR comparing BM25 and Contriever. We use the same seed set and corpus as Table 2. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. The top two subsections consider zero-shot synthesis and bottom two considers 3-shot Retr-ICL variant. We bold the best result in each subsection. Contriever numbers are reproduced from Table 6.
So far, a single dense retriever (Contriever) has been used for the content sourcing step by using a bi-encoder approach (Lee et al., 2019; Chen et al., 2017). We embed both the input in-context covariate and each corpus document, and then rank results based on cosine similarity. In §5, we retrieved $k=500$ documents for each in-context example and after filtering, randomly sampled among these to produce a grounded set of documents on which we apply our task inversion strategy Retr-ICL.
In this section we explore how changing the retrieval model affects the content sourcing stage and its downstream effects. Table 11 compares the DeBERTa-v3L accuracy of Contriever and BM25. Keeping other parts of the process the same, we switch Contriever to BM25 Okapi (Robertson and Zaragoza, 2009), a popular sparse retrieval method which performs a lexical match based on inverse term frequencies. Unlike dense retrieval models, BM25 outputs a score which is an unbounded positive number. Thus we are unable to use meaningful thresholds to bound the similarity in our Retr-ICL approach. Instead, we construct the Retr-ICL in-context set using the top-2 retrieved contexts for each ICL example and without applying the filter.
We observe in Table 11 that a powerful student model trained on the dense-retrieved document set consistently outperforms the sparse-retrieved documents from BM25. In Table 10, it is also observed that the intrinsic diversity from BM25-retrieved documents is also not as good as that from Contriever.
Dense retrievers like Contriever perform a semantic match between the query and document, whereas BM25 performs only a lexical match with no understanding of semantics. Picking semantically similar information is more important in SynthesizRR since we include a task inversion step, which can change the lexical structure of the text without changing its semantics significantly. Thus, we want contexts which are semantically related to Gold data, to which we can apply stylistic or formatting transformations using a task-inversion prompt to bring it closer to Gold. We might conclude that dense retrieval models are more suitable for SynthesizRR.
Appendix C Influence of retrieval corpus on domain shift

### v3

@@ Appendix B Incorporating feedback from distilled student models
RQ: Why does SynthesizRR improve classification dataset synthesis? In this section we take a closer look at the generated classification dataset and how it affects the training dynamics of student models during distillation.
Aside from the final accuracy, we also consider label preservation accuracy, which is obtained from an “oracle” model for the task. We construct this oracle from Gold data by running a grid-search over DeBERTa-v3-Large hyperparams (Appendix J), splitting 80% of the Gold train set for fine-tuning and 20% for validation. Then, we measure the fraction of synthetic examples which the oracle classifies to belong to the prompted target class. This indicates the adherence of the generated example to the class it should belong to, as per the prompt.
@@
We would expect that better label preservation means a higher-fidelity training dataset. However, Table 8 shows that FewGen datasets have very high label preservation in spite of their lower test performance. Especially on multiclass tasks (AG., ToI, Cat.), FewGen shows the highest label preservation (exceeding Gold) but this does not translate into improved student performance.
@@
To understand this, we conduct a deeper analysis of the student training dynamics on multiclass datasets. We train a DistilBERT student for 6 epochs and plot the corresponding data-maps Swayamdipta et al. (2020). For binary tasks, the data-maps for SynthesizRR matched both FewGen and Gold, but the data maps from multi-class differed greatly. Figure 5 illustrates this difference using the Category task maps. From Figure 5 it is clear that FewGen generations tend to cluster around easy-to-learn examples (high confidence and low variability), whereas SynthesizRR contains more ambiguous examples (high variability) which Swayamdipta et al. (2020) demonstrate is essential to learning the nuances between classes.
RQ: Can we improve distillation performance by leveraging student feedback from data-maps?
Swayamdipta et al. (2020) use data-maps to filter out easy to-learn examples (top-left, red) and potentially mislabelled examples (bottom-left, blue) and obtain superior accuracy on human-generated datasets. We attempt to apply this same technique to the synthetic datasets generated by SynthesizRR and FewGen.
Concretely, we filter out the least ambiguous examples (bottom 17% variability) and retrain the DistilBERT student model on the smaller, filtered dataset. In Table 9 we find that FewGen performance degrades, whereas SynthesizRR improves (giving us new best performances on multi-class despite using only 83% of rows). We conclude that SynthesizRR generates more ambiguous examples, and this helps establish better class-separability in multi-class data sets.
@@
SynzthRR $87.56$ 72.8 $74.79$ $69.41$ 90.7 $99.32$
SynzthRR $87.36$ $65.85$ $73.24$ $73.24$ $77.35$ 99.7
Table 8: Few-shot label-preservation accuracy ($\uparrow$) using tuned oracle DeBERTa-v3L model. Gold row is accuracy on 20% validation split. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
@@
Table 9: Test Accuracy ()($\uparrow$)( ↑ ) after keeping 83% most-ambiguous examples. We report improvements compared to Table 6. Notation: *32-shot; 3-shot RetrICL; 32-shot Non-RetrICL.
Appendix C Bootstrapping with a synthetic seed set
A core assumption in SynthesizRR has been the existence of a small seed set of human-written $(x,y)$ pairs for the task. This seed set is critical as it serves a dual purpose: it is used as the set of the retrieval queries, and as in-context learning examples to guide the teacher LLM’s next-token distribution in the task inversion step.
In this section we consider how we can synthesize such a seed set for low-resource settings. Our core assumption is that the seed set is small (100/class for binary tasks and 50/class for multiclass tasks). Thus using FewGen with top-$p=0.9$ and temperature $=0.95$ and three in-context examples, we attempt to generate a diverse seed set with minimal repetitions. This bootstrapping approach makes SynthesizRR tractable when very little human data is available (just 5-15 examples per class) or no human data is available.
Concretely, we compare three paradigms:
  1. 1.
    True zero-shot: when we have no human data we utilize zero-shot generation to bootstrap the seed set.
  2. 2.
    Low-resource: Here, we assume we have a small number of human-written examples, e.g. 5 examples per class. This is presumed insufficient to be used as the seed set directly, but we can use it as in-context examples to guide the FewGen generator to bootstrap a realistic seed set.
  3. 3.
    Sufficient: We do not synthesize the seed set. This is the SynthesizRR paradigm we have explored in previous sections, wherein we have 50-100 Gold examples per class in our seed set.
As mentioned in §4, the true zero-shot paradigm makes strong assumptions that are often unnecessarily restrictive. In practice, it is typically feasible to obtain a small amount of human-written examples (low-resource or sufficient seed), while obtaining several thousand human-written examples is still challenging.
Gold RetrICL AG. Hyp. ToI Cat. Hum. Pol.
data ($N$) shots ($8$K) ($2$K) ($8$K) ($8$K) ($2$K) ($4$K)
Gold
All - 91.0 93.2 82.5 81.5 93.1 95.3
True Zero-shot (0-shot FewGen seed)
None 0-shot 66.6 68.0 60.5 60.4 76.9 76.4
None 3-shot 60.0 72.3 62.5 61.7 72.3 85.4
Low-Resource ($\binom{N}{3}$-shot FewGen seed)
5/class 0-shot 79.9 71.7 68.1 63.4 81.3 81.3
5/class 3-shot 77.7 66.8 68.9 58.8 86.4 86.5
15/class 0-shot 78.5 72.9 69.3 65.7 77.4 84.0
15/class 3-shot 76.1 72.6 71.6 63.5 82.5 73.8
Sufficient (Gold Seed)
Full seed 0-shot 83.5 69.8 74.5 68.9 82.5 84.7
Full seed 3-shot 83.0 78.5 73.3 72.4 90.2 91.0
Table 10: Test accuracy after distilling a DeBERTa-v3L student on a dataset generated by SynthesizRR RetrICL variant. We use the same corpus as Table 2, but vary the seed set. LLaMa-2 Chat 13B is used as the teacher LLM. We train 5 student models and report the mean accuracy, rerunning all 5 in case of std $\geq$ 6.0. “’Full” seed implies 100 Gold examples per class for binary and 50 per class for multiclass tasks. We bold the best result in each paradigm.
The results of running SynthesizRR RetrICL using synthetic seed data is shown in Table 10. As a general trend, adding more human-written examples leads to better performance. Unsurprisingly, the best results are in the Sufficient paradigm, where we use 50-100 Gold examples as both retrieval queries and the the RetrICL set. True Zero-shot results (without any human input) are considerably worse. Surprisingly, however, we are able to get good distillation accuracy with just 5 examples per class rather than the full 50-100 per class, which indicates that SynthesizRR might be usable in low-resource settings where human annotated data is scarce.
In certain cases of the low-resource paradigm, we observe that the performance drops significantly from 0-shot RetrICL to 3-shot RetrICL. We attribute this to the fact that, even with 5-15 Gold in-context examples, the FewGen-generated seed set might not be reflective of the true Gold examples (this behavior is reflected in the low MAUVE scores in Table 5). Thus, by conditioning on incorrect synthetic examples during RetrICL, we shift the next-token distribution away from the true distribution.
In conclusion, using FewGen to bootstrap a seed set can be a viable approach to using SynthesizRR in low-resource settings where there is not enough Gold task-data.
Appendix D Influence of corpus on domain shift